1. What is the value (in V) of Eocell for the following reaction?

Co2+ (aq) + Be (s) → Co (s) + Be2+ (aq)

2.What is Eocell (in V) for a redox reaction where one electron is transferred with an equilibrium constant (K) of 1.44 x 10-14?

3.Consider the following reaction:

Cu2+ (aq) + Pb (s) → Cu (s) + Pb2+ (aq)

What will be Ecell for this reaction (in V) when [Cu2+] = 0.500 M and [Pb2+] = 0.0350 M, and the temperature is 298 K?
Answer #1

1)

Consider reactions taking place into the cell.

Reaction at anode: Be (s) Be 2+ (aq) + 2 e -

Reaction at cathode : Co 2+ (aq) + 2 e - Co (s)

Overall cell reaction :Be (s) + Co 2+ (aq)   Be 2+ (aq) + Co (s)

The standard emf of the cell is E 0 cell = E 0 cathode - E 0 anode

E 0 cell = - 0.282 V - ( - 1.968 V )

E 0 cell = 1.686 V

ANSWER : E 0 cell = 1.686 V

2)

We have relation, E 0 cell = 2.303 R T / n F log K c

E 0 cell = ( 2.303 8.314 298 / 1 96487 ) log  1.44 10-14

E 0 cell = - 0.8185 V

ANSWER : E 0 cell = - 0.818 V

3)

Consider reactions taking place into the cell

At anode: Pb (s) Pb 2+(aq) + 2 e -

At cathode: 2 Cu 2+ (aq) + 2 e -    2 Cu(s)

Overall reaction : Cu 2+ (aq) + Pb (s) Cu (s) + Pb 2+ (aq)

No. of electrons transferred in above reaction = 2

Standard emf of the cell is calculated as, E 0 cell = E 0 cathode - E 0 anode

E 0 cell = 0.339 V - ( - 0.126 V )

E 0 cell = 0.465 V

We have Nerns't Equation, E cell = E 0 cell - [2.303 R T / n F ] log Q

At 25 0 C, 2.303 R T/ F = 0.0591 Therefore, we can write

E cell = E 0 cell - 0.0591 / n log [Pb 2+  ] / [ Cu 2+ ]

= 0.465 - 0.0592 / 2 log 0.0350 / 0.500

= 0.465 - ( - 0.0342 )

= 0.464 + 0.0342

= 0.499 V

ANSWER : E cell = 0.499 V